# Simulating Disky Broad Line Region Reverberation

### v1

@@ 1 Introduction
The optical spectra of active galactic nuclei (AGN) and quasars often show broad emission lines khachikian74 that have widths on the order of km s-1 (e.g., peterson93). The origin of these broad emission lines is the gaseous region correspondingly called the Broad-Line Region (BLR), which is photoionized by the high-energy continuum coming from the central engine of the AGN. Observations of the emission lines show the same variability pattern as the observed optical and UV continuum with a smaller amplitude and a time delay. This behavior is known as reverberation of the BLR. Reverberation mapping blandford82 can be used to probe the structure and kinematics of the BLR, and has obtained constraints on the extent of the BLR and provides most of the information currently known on BLR sizes. Only a few BLR geometries have been successfully spatially resolved using interferometric methods (e.g., gravity20), which further confirm the results of reverberation mapping.
@@
Many intensive reverberation mapping campaigns have been carried out, starting with efforts in the 1990s (see clavel91; korista95) and progressing to more modern efforts, such as the LAMP campaign (see bentz09), AGN STORM and AGN STORM 2 (see derosa15; kara21), and the Sloan Digital Sky Surveys (see eisenstein11; blanton17; shen15). As a result of such campaigns, a number of behaviors have been identified observationally in the reverberation signatures of the BLR. A particular behavior that is vital for obtaining black hole masses in these campaigns in that the BLR is “virialized” (peterson99; grier13; u22), or that the gas in the BLR obeys the virial theorem (peterson99; peterson00). Another behavior observed is the “radius-luminosity” relation – a correlation between time lag and the continuum luminosity (e.g., kaspi00; bentz13).
@@
Other studies have shown that there is an anti-correlation between the luminosity of the AGN and the broad emission line widths, notably when examining the H line (e.g., dexter19). This effect, known as ‘breathing’, can produce a narrower line profile by boosting the core’s low velocity flux compared to the higher velocity wings of the profile. korista04 and goad14 attribute this behavior to the higher responsivity of H farther out in the BLR. Therefore, when the continuum changes, the response of the core of the line is more pronounced, and produces this effect (e.g., when the continuum strength increases the core of the line responds more strongly than the wings and the line appears narrower). barth15 and park12 have found that breathing, previously thought to appear over dynamical timescales of the BLR (of order several years, see obrien95; korista04; cackett06; bentz07), can appear on timescales of the order of days to weeks following changes in the continuum. Although barth15 and park12 found that the broad H lines in most of the 15 objects they studied exhibited breathing behavior, it is important to note that “anti-breathing” has also been observed: a study by guo14 found that the majority of their sample of 60 objects exhibited “anti-breathing” behavior. zhang15 also examined an object in more depth that exhibits anti-breathing behavior in some of the broad emission lines, and wang20 find that different broad emission lines may show breathing or anti-breathing in the same object. The centroid velocity shift of these lines has been found to vary on similar timescales as breathing barth15. This effect has been termed “jitter”, or “radial velocity jitter”. However, velocity jitter has also been observed to vary over longer timescales (see sergeev07). In some cases jitter has been correlated with changes in the ionizing continuum or luminosity of the central engine, which may indicate that it may be a result of the BLR reverberation.
@@
The role of radial velocity jitter is also vital in spectroscopic searches for binary SMBHs (e.g., popovic12; dotti12; bogdanovic15). Under the assumption that one SMBH in the binary is active and has a BLR, the broad emission lines will oscillate in velocity over an orbital period and can reveal sub-parsec binaries that cannot otherwise be telescopically resolved. The expected orbital periods of the binary SMBHs that can be found using radial velocity variations are on the order of several decades to centuries (e.g., eracleous12). Most of the promising spectroscopic times series currently available span under 30 years and are sparsely sampled (doan20; mohammed25). Radial velocity jitter that is a result of the reverberation of the BLR can create apparent variations that can be mistaken for the signature of orbital motion. Therefore, radial velocity jitter of the emission line profiles must be characterized and understood in order to appropriately deal with this effect in SMBH binary searches.
@@
In this paper we present the design of a software pipeline that can simulate reverberation of a disk-like BLR, which is achieved by shining a model for the light curve from an AGN central engine on the BLR, and then calculating the subsequent time-dependent emission line profiles. This pipeline takes advantage of published physical models to specify the geometry, kinematics, emissivity laws, and other features of the BLR, and incorporates special and general relativistic effects, which capture asymmetries in the line profiles even when the BLR is axisymmetric. We focus specifically on a disk-like BLR in this work because this model is well developed theoretically and has substantial observational support. We justify the use of this model further in Section 2. Previous studies (e.g., welsh91; robinson95) have not assumed a single model of the BLR, but instead used more general reverberation mapping techniques (such as transfer functions) to get better constraints on the geometry and kinematics using multiple proposed models. There are also a number of forward-modeling codes that are very sophisticated and can adopt a number of geometries and velocity fields in order to produce velocity response functions for a variety of applications; see pancoast11, li13, and rosborough24. Such forward-modeling codes focus on obtaining the mass of the SMBH from the constraints on the geometry of BLR, and seek not to use the virial factor, , that is used in traditional reverberation mapping.
@@
Our project and code are specifically targeting the disk-like geometry of the BLR and explore how the broad emission line profiles may change as a result of this geometry when it reverberates from the central ionizing continuum. Our code uses analytic prescriptions instead of photoionization grids to save computational time and incorporates the effects of special and general relativity in the calculation of the line profiles, which were shown to be important (tremaine14). Thus, our pipeline allows us to quickly explore a very large parameter space for the model. A more thorough exploration of a specific range of parameter combinations may have to employ more sophisticated codes like those cited above, after our code has identified that parameter space. We aim to use this pipeline in forthcoming papers to study jitter and breathing of the broad Balmer lines. Our broader goals are to exploit breathing and jitter to develop tests for the disk-like BLR model and to explore the possibility of developing a viable jitter prescription for other applications. The pipeline was developed and tested for one particular BLR geometry, but is modular and can be adapted to apply to other geometries and velocity fields of the BLR as needed.
@@
This paper is the first in a series and describes the development of the machinery for this project–future papers will discuss science results. In Section 2, we introduce the model we are using, as well as the important assumptions and features that are included within the model. In Section 3, we give a brief overview of the main modules of the pipeline and the process necessary to create the simulated emission line profiles. In Section 4, we detail the exact function of each module. In Section 5, we discuss the current stage of the pipeline and describe end-to-end tests. We close with a summary and a brief outline of future work at the end of Section 5.
2 Description of the Model
2.1 Geometry and Structure of the Broad-Line Region
@@
Our goal in this version of the pipeline is to simulate the observational signature of a disk-like BLR and compute the variability of the moments of the resulting line profiles. We are assuming that the BLR has the form of a Keplerian disk that is just an extension of the hot inner accretion disk that produces the ionizing continuum. We incorporate analytic prescriptions of physical effects, such as radiative transfer in the base of the disk wind, relativistic effects, and emissivities intrinsic to the BLR to understand the changes on the line profiles.
@@
A disk-like BLR is a well-developed physical model and is consistent with the observed behaviors noted in the previous section—thus this pipeline can be used to test the model further. The BLR in this model is a continuous medium that is consistent with a disk and/or wind, and reverberation mapping campaigns often show the signature of a reverberating disk (e.g., horne21; bentz23). The VLT Interferometer has also resolved the BLR in 3C 273 and revealed that the BLR looks like a disk (e.g., gravity20). Several authors have considered models for a disk-like BLR, with some of the earlier models coming from emmering92, the series by collin-souffrin89, murray95, and elitzur14. More modern and sophisticated models are based on the same idea and include those of waters16, matthews20, baskin18, and naddaf22. The disk-like model for this project combines and utilizes the ideas from the above papers.
@@
The continuum originates in a central source much smaller than the BLR—i.e., a point source dumont90a.
The BLR is in the equatorial plane and makes up the outer skin of the accretion disk. The emission layer may be accelerating in the vertical direction to form a wind.
The emission lines respond linearly (to first order) to variations of the ionizing continuum (see Section 5.1 for further discussion).
@@
The light-travel time across the BLR is the most important time scale—the local gas response is effectively instantaneous with a response time of seconds (peterson93), and the light-travel time is much shorter than the dynamical time scale for the BLR.
@@
2.2 Calculation of the Line Profiles
We follow the approach described in flohic12 to calculate the broad emission line profiles from a disk-like BLR. This formalism, represented in Eqn. 1, was obtained by combining the methods described in chen89 and chen89b with those of murray95. We adopt the geometry and coordinate system in Figure 1 of chen89 and assume that the line emitting region is an annulus defined by an inner and an outer radius. We adopt a convention in which the unprimed coordinates () refer to the frame of the disk. We numerically calculate the flux density of the line profile as seen by an observer at infinity through
@@ 2.3 Emissivity Law(s)
collin-souffrin89 introduced both single and broken power-law approximations for in order to describe the results of their photoionization calculations. dumont90b have shown that the emitted flux of the H line drops as as a result of the geometric dilution of illumination. Several papers, such as eracleous94, eracleous03, storchi-bergmann17, and ward24, have found that a power-law emissivity leads to good agreement between models and observed double-peaked line profiles. These studies have also found a similar agreement with the implementation of a broken power-law. Thus, the emissivity can be described by:
@@
for a broken power-law, where is the break radius, and and are the indices for the emissivity in their respective radial regions. Figure 1 shows an example of a disk with an emissivity of the form . A separate, empirical law is inspired by reverberation mapping campaigns and can be parameterized as a bell-shaped function that peaks at some radius and declines in either direction (e.g., horne21). This law can be represented as
where is the radial peak of the Gaussian, and . This bell-shaped function can subsequently be combined with a radial or broken power-law (see sottocorno25, for examples) in order to reflect the empirical relation between the BLR characteristic radius and central engine luminosity, as discussed in Section 1.
Non-axisymmetric perturbations can be added to the disk, which can explain the shapes and variability of Balmer lines. Spiral arms can form in the disk as a result of gravitational instabilities within a self-gravitating disk or external perturbations (e.g., adams89; livio91; chakrabarti93a).
Previous works have found that the inclusion of a static or precessing spiral arm or bright spot on the surface of the disk leads to good fits of broad emission line profiles for single objects and samples of objects, (e.g., storchi-bergmann03; gezari07; lewis10; storchi-bergmann17; ward24; ward25; rigamonti25). Thus we use the spiral arm prescription of gilbert99 and storchi-bergmann03:
@@
where is a fiducial radius (typically the inner radius of the disk), is the pitch angle of the spiral arm, and sets the orientation of the pattern. An example of a spiral arm emissivity is given in Figure 2. The values of the spiral arm parameters are constrained by fits to observed line profiles (see schimoia12; schimoia17; storchi-bergmann17; ward24; ward25, references therein). We have narrowed the parameter range to , , , and according to these studies.
@@
The final component of the overall emissivity of the disk is the directional escape probability. In the models of murray95, the broad lines are produced at the base of a disk wind and must travel through the accelerating emission layer to escape. We use the following expression for the escape probability (hamann93; murray98)
@@
In order to calculate the position-dependent optical depth , we utilize the following expression from flohic12 (see also murray98; chajet13)
@@
where is the total optical depth of the surface emission layer perpendicular to the disk plane at , describes the dependence of density in the emission layer on radius, and represents the analytic prescription of the projected strain tensor of a disk wind from murray95 (see Eqn. 8 in flohic12, for the full expression for and its dependencies on the inclination angle between the observer and the axis of the disk, ). An illustrative example of how the escape probability affects the emissivity of the disk can be seen in Figure 3.
2.4 Models for the Light Curve

### v2

@@ I Introduction
The optical spectra of active galactic nuclei (AGN) and quasars often show broad emission lines Khachikian and Weedman (1974) that have widths on the order of km s-1 (e.g., Peterson, 1993). The origin of these broad emission lines is the gaseous region correspondingly called the Broad-Line Region (BLR), which is photoionized by the high-energy continuum coming from the central engine of the AGN. Observations of the emission lines show the same variability pattern as the observed optical and UV continuum with a smaller amplitude and a time delay. This behavior is known as reverberation of the BLR. Reverberation mapping Blandford and McKee (1982) can be used to probe the structure and kinematics of the BLR, and has obtained constraints on the extent of the BLR and provides most of the information currently known on BLR sizes. Only a few BLR geometries have been successfully spatially resolved using interferometric methods (e.g., GRAVITY Collaboration et al., 2020), which further confirm the results of reverberation mapping.
@@
Many intensive reverberation mapping campaigns have been carried out, starting with efforts in the 1990s (see Clavel et al., 1991; Korista et al., 1995) and progressing to more modern efforts, such as the LAMP campaign (see Bentz et al., 2009), AGN STORM and AGN STORM 2 (see De Rosa et al., 2015; Kara et al., 2021), the Sloan Digital Sky Surveys (see Eisenstein et al., 2011; Blanton et al., 2017; Shen et al., 2015), the Supermassive Black Holes with High Accretion Rates in Active Galactic Nuclei program (see Du et al., 2014), and the Seoul National University AGN Monitoring Project (see Woo et al., 2019). As a result of such campaigns, a number of behaviors have been identified observationally in the reverberation signatures of the BLR. A particular behavior that is vital for obtaining black hole masses in these campaigns in that the BLR is “virialized” (Peterson and Wandel, 1999; Grier et al., 2013; U et al., 2022), or that the gas in the BLR obeys the virial theorem (Peterson and Wandel, 1999, 2000). Another behavior observed is the “radius-luminosity” relation – a correlation between time lag and the continuum luminosity (e.g., Kaspi et al., 2000; Bentz et al., 2013).
@@
Other studies have shown that there is an anti-correlation between the luminosity of the AGN and the broad emission line widths, notably when examining the H line (e.g., Dexter et al., 2019). This effect, known as ‘breathing’, can produce a narrower line profile by boosting the core’s low velocity flux compared to the higher velocity wings of the profile. Korista and Goad (2004) and Goad and Korista (2014) attribute this behavior to the higher responsivity of H farther out in the BLR. Therefore, when the continuum changes, the response of the core of the line is more pronounced, and produces this effect (e.g., when the continuum strength increases the core of the line responds more strongly than the wings and the line appears narrower). Barth et al. (2015) and Park et al. (2012) have found that breathing, previously thought to appear over dynamical timescales of the BLR (of order several years, see O’Brien et al., 1995; Korista and Goad, 2004; Cackett and Horne, 2006; Bentz et al., 2007), can appear on timescales of the order of days to weeks following changes in the continuum. Although Barth et al. (2015) and Park et al. (2012) found that the broad H lines in most of the 15 objects they studied exhibited breathing behavior, it is important to note that “anti-breathing” has also been observed: a study by Guo and Gu (2014) found that the majority of their sample of 60 objects exhibited “anti-breathing” behavior. Zhang (2015) also examined an object in more depth that exhibits anti-breathing behavior in some of the broad emission lines, and Wang et al. (2020) find that different broad emission lines may show breathing or anti-breathing in the same object. The centroid velocity shift of these lines has been found to vary on similar timescales as breathing Barth et al. (2015). This effect has been termed “jitter”, or “radial velocity jitter”. However, velocity jitter has also been observed to vary over longer timescales (see Sergeev et al., 2007). In some cases jitter has been correlated with changes in the ionizing continuum or luminosity of the central engine, which may indicate that it may be a result of the BLR reverberation.
@@
The role of radial velocity jitter is also vital in spectroscopic searches for binary SMBHs (e.g., Popović, 2012; Dotti et al., 2012; Bogdanović, 2015). Under the assumption that one SMBH in the binary is active and has a BLR, the broad emission lines will oscillate in velocity over an orbital period and can reveal sub-parsec binaries that cannot otherwise be telescopically resolved. The expected orbital periods of the binary SMBHs that can be found using radial velocity variations are on the order of several decades to centuries (e.g., Eracleous et al., 2012). Most of the promising spectroscopic times series currently available span under 30 years and are sparsely sampled (Doan et al., 2020; Mohammed et al., 2025). Radial velocity jitter that is a result of the reverberation of the BLR can create apparent variations that can be mistaken for the signature of orbital motion. Therefore, radial velocity jitter of the emission line profiles must be characterized and understood in order to appropriately deal with this effect in SMBH binary searches.
@@
In this paper we present the design of a software pipeline that can simulate reverberation of a disk-like BLR, which is achieved by shining a model for the light curve from an AGN central engine on the BLR, and then calculating the subsequent time-dependent emission line profiles. This pipeline takes advantage of published physical models to specify the geometry, kinematics, emissivity laws, and other features of the BLR, and incorporates special and general relativistic effects, which capture asymmetries in the line profiles even when the BLR is axisymmetric. We focus specifically on a disk-like BLR in this work because this model is well developed theoretically and has substantial observational support. We justify the use of this model further in Section II. Previous studies (e.g., Welsh and Horne, 1991; Robinson, 1995) have not assumed a single model of the BLR, but instead used more general reverberation mapping techniques (such as transfer functions) to get better constraints on the geometry and kinematics using multiple proposed models. There are also a number of forward-modeling codes that are very sophisticated and can adopt a number of geometries and velocity fields in order to produce velocity response functions for a variety of applications; see Pancoast et al. (2011), Li et al. (2013), and Rosborough et al. (2024). Such forward-modeling codes focus on obtaining the mass of the SMBH from the constraints on the geometry of BLR, and seek not to use the virial factor, , that is used in traditional reverberation mapping.
@@
Our project and code are specifically targeting the disk-like geometry of the BLR and explore how the broad emission line profiles may change as a result of this geometry when it reverberates from the central ionizing continuum. Our code uses analytic prescriptions instead of photoionization grids to save computational time and incorporates the effects of special and general relativity in the calculation of the line profiles, which were shown to be important (Tremaine et al., 2014). Thus, our pipeline allows us to quickly explore a very large parameter space for the model. A more thorough exploration of a specific range of parameter combinations may have to employ more sophisticated codes like those cited above, after our code has identified that parameter space. We aim to use this pipeline in forthcoming papers to study jitter and breathing of the broad Balmer lines. Our broader goals are to exploit breathing and jitter to develop tests for the disk-like BLR model and to explore the possibility of developing a viable jitter prescription for other applications. The pipeline was developed and tested for one particular BLR geometry, but is modular and can be adapted to apply to other geometries and velocity fields of the BLR as needed.
@@
This paper is the first in a series and describes the development of the machinery for this project–future papers will discuss science results. In Section II, we introduce the model we are using, as well as the important assumptions and features that are included within the model. In Section III, we give a brief overview of the main modules of the pipeline and the process necessary to create the simulated emission line profiles. In Section IV, we detail the exact function of each module. In Section V, we discuss the current stage of the pipeline and describe end-to-end tests. We close with a summary and a brief outline of future work at the end of Section V.
II Description of the Model
II.1 Geometry and Structure of the Broad-Line Region
@@
Our goal in this version of the pipeline is to simulate the observational signature of a disk-like BLR and compute the variability of the moments of the resulting line profiles. We are assuming that the BLR has the form of a Keplerian disk that is just an extension of the hot inner accretion disk that produces the ionizing continuum. In this pipeline, we make no assumptions regarding the hot inner accretion disk. The hot inner accretion disk can be described by any model that allows it to emit ionizing radiation that can reach the BLR either through direct or indirect illumination. We incorporate analytic prescriptions of physical effects, such as radiative transfer in the base of the disk wind, relativistic effects, and emissivities intrinsic to the BLR to understand the changes on the line profiles.
@@
A disk-like BLR is a well-developed physical model and is consistent with the observed behaviors noted in the previous section—thus this pipeline can be used to test the model further. The BLR in this model is a continuous medium that is consistent with a disk and/or wind, and reverberation mapping campaigns often show the signature of a reverberating disk (e.g., Horne et al., 2021; Bentz et al., 2023). The VLT Interferometer has also resolved the BLR in 3C 273 and revealed that the BLR looks like a disk (e.g., GRAVITY Collaboration et al., 2020). Several authors have considered models for a disk-like BLR, with some of the earlier models coming from Emmering et al. (1992), the series by Collin-Souffrin and Dumont (1989), Murray et al. (1995), and Elitzur et al. (2014). More modern and sophisticated models are based on the same idea and include those of Waters et al. (2016), Matthews et al. (2020), Baskin and Laor (2018), and Naddaf and Czerny (2022). The disk-like model for this project combines and utilizes the ideas from the above papers. It is possible for users of this pipeline to adopt other disk geometries; in such a case the user would have to prescribe the time sequences of emissivity maps accordingly through the apparent emissivity generator and brightness map modules (discussed in Section II.3).
@@
The continuum originates in a central source much smaller than the BLR—i.e., a point source Dumont and Collin-Souffrin (1990a).
The BLR is in the equatorial plane and the line-emitting region makes up the “skin” (i.e., the surface emission layer) of the outer disk. The emission layer may be accelerating in the vertical direction to form a wind.
The emission lines respond linearly (to first order) to variations of the ionizing continuum (see Section V.1 for further discussion).
@@
The light-travel time across the BLR is the most important time scale—the local gas response is effectively instantaneous with a response time of seconds (Peterson, 1993), and the light-travel time is much shorter than the dynamical time scale for the BLR.
@@
II.2 Calculation of the Line Profiles
We follow the approach described in Flohic et al. (2012) to calculate the broad emission line profiles from a disk-like BLR. This formalism, represented in Eqn. 1, was obtained by combining the methods described in Chen et al. (1989) and Chen and Halpern (1989) with those of Murray et al. (1995). We adopt the geometry and coordinate system in Figure 1 of Chen et al. (1989) and assume that the line emitting region is an annulus defined by an inner and an outer radius. We adopt a convention in which the unprimed coordinates () refer to the frame of the disk. We numerically calculate the flux density of the line profile as seen by an observer at infinity through
@@ II.3 Emissivity Law(s)
Collin-Souffrin and Dumont (1989) introduced both single and broken power-law approximations for in order to describe the results of their photoionization calculations. The power law is not a direct prediction of their physical model but a convenient parametrization for their numerical calculations, and Dumont and Collin-Souffrin (1990b) have shown that the emitted flux of the H line drops as as a result of the geometric dilution of illumination. The use of parametric emissivity laws allows users of the pipeline to choose the line emissivity according to their goals, whether that be from numerical or analytic calculations, or empirical results. Several papers, such as Eracleous and Halpern (1994), Eracleous and Halpern (2003), Storchi-Bergmann et al. (2017), and Ward et al. (2024), have found that a power-law emissivity leads to good agreement between models and observed double-peaked line profiles. These studies have also found a similar agreement with the implementation of a broken power-law. Thus, the emissivity can be described by:
@@
for a broken power-law, where is the break radius, and and are the indices for the emissivity in their respective radial regions. Figure 1 shows an example of a disk with an emissivity of the form . A separate, empirical law is inspired by reverberation mapping campaigns and can be parameterized as a bell-shaped function that peaks at some radius and declines in either direction (e.g., Horne et al., 2021). This law can be represented as
where is the radial peak of the Gaussian, and . This bell-shaped function can subsequently be combined with a radial or broken power-law (see Sottocorno et al., 2025, for examples) in order to reflect the empirical relation between the BLR characteristic radius and central engine luminosity, as discussed in Section I.
Non-axisymmetric perturbations can be added to the disk, which can explain the shapes and variability of Balmer lines. Spiral arms can form in the disk as a result of gravitational instabilities within a self-gravitating disk or external perturbations (e.g., Adams et al., 1989; Livio and Spruit, 1991; Chakrabarti and Wiita, 1993).
Previous works have found that the inclusion of a static or precessing spiral arm or bright spot on the surface of the disk leads to good fits of broad emission line profiles for single objects and samples of objects, (e.g., Storchi-Bergmann et al., 2003; Gezari et al., 2007; Lewis et al., 2010; Storchi-Bergmann et al., 2017; Ward et al., 2024, 2025; Rigamonti et al., 2025). Thus we use the spiral arm prescription of Gilbert et al. (1999) and Storchi-Bergmann et al. (2003):
@@
where is a fiducial radius (typically the inner radius of the disk), is the pitch angle of the spiral arm, and sets the orientation of the pattern. An example of a spiral arm emissivity is given in Figure 2. The values of the spiral arm parameters are constrained by fits to observed line profiles (see Schimoia et al., 2012, 2017; Storchi-Bergmann et al., 2017; Ward et al., 2024, 2025, references therein). We have narrowed the parameter range to , , , and according to these studies.
@@
The final component of the overall emissivity of the disk is the directional escape probability. In the models of Murray et al. (1995), the broad lines are produced at the base of a disk wind and must travel through the accelerating emission layer to escape. We use the following expression for the escape probability (Hamann et al., 1993; Murray and Chiang, 1998)
@@
In order to calculate the position-dependent optical depth , we utilize the following expression from Flohic et al. (2012) (see also Murray and Chiang, 1998; Chajet and Hall, 2013)
@@
where is the total optical depth of the surface emission layer perpendicular to the disk plane at , describes the dependence of density in the emission layer on radius, and represents the analytic prescription of the projected strain tensor of a disk wind from Murray et al. (1995) (see Eqn. 8 in Flohic et al., 2012, for the full expression for and its dependencies on the inclination angle between the observer and the axis of the disk, ). An illustrative example of how the escape probability affects the emissivity of the disk can be seen in Figure 3.
Through the generation of these maps, we dictate how the illumination sets the emissivity of the BLR. Users can modify the relevant modules to produce maps that encode geometric or emissivity information according to their own goals. For example, one can adopt an non-axisymmetric illuminaiton pattern for the BLR and compute the maps accordingly.
II.4 Models for the Light Curve